# DAWN: Designing Distributed Agents in a Worldwide Network

### v2

@@ Abstract
The rapid evolution of Large Language Models (LLMs) has transformed them from basic conversational tools into sophisticated entities capable of complex reasoning and decision-making. These advancements have led to the development of specialized LLM-based agents designed for diverse tasks such as coding and web browsing. As these agents become more capable, the need for a robust framework that facilitates global communication and collaboration among them towards advanced objectives has become increasingly critical. Distributed Agents in a Worldwide Network (DAWN) addresses this need by offering a versatile framework that integrates LLM-based agents with traditional software systems, enabling the creation of agentic applications suited for a wide range of use cases. DAWN enables distributed agents worldwide to register and be easily discovered through Gateway Agents. Collaborations among these agents are coordinated by a Principal Agent equipped with reasoning strategies. DAWN offers three operational modes: No-LLM Mode for deterministic tasks, Copilot for augmented decision-making, and LLM Agent for autonomous operations. Additionally, DAWN ensures the safety and security of agent collaborations globally through a dedicated safety, security, and compliance layer, protecting the network against attackers and adhering to stringent security and compliance standards. These features make DAWN a robust network for deploying agent-based applications across various industries.
Index Terms:
Large Language Model, Agent, Multi-agent system
@@ I Introduction
Rapid advancements in large language models (LLMs)[1, 2, 3] have have sparked significant interest in LLM-based agents[4] which have the potential to free humans from repetitive tasks and significantly enhance productivity. New agents are being developed and their performance is improving on a regular basis. There are agents for many tasks such as coding[5, 6], web browsing[7], gaming, and many more.[8, 9]. Furthermore, multi-agent systems[10, 11, 12, 13, 14, 15, 16] have been proposed to leverage the collective intelligence and specialized profiles and skills of multiple agents. However, three major challenges arise that hinder their potential for broader, more impactful applications.
@@
The first challenge is how to discover and connect to agents distributed globally and accomplish complex tasks. Existing frameworks, such as LangGraph[14] and AutoGen[10] have bounded scope as they are typically designed to manage a limited number of agents. This constraint arises from the need to manually hard-code the names, URLs, and description of other agents into the prompts, which becomes increasingly difficult as the network of agents grows. More agents allow for a broader selection covering functional and non-functional capabilities. Additionally, integrating heterogeneous agents from different ecosystems requires developers to invest effort in developing wrappers and packages specific to each new framework as they evolve.
@@
The second challenge is how to incorporate determinism into agentic applications, particularly when tasks require both flexible decision-making and predictable, reliable outcomes. While LLM-based agents excel in dynamic environments that exhibit flexibility, they falter in tasks requiring strict adherence to set execution graphs, predefined rules, and outcomes. Empirical evidence shows that manually designing a workflow rather than allowing agents to autonomously make decisions leads to more robust performance and cuts down on computational cost. Balancing the often orthogonal needs for autonomy and determinism is essential for the broader adoption of agent-based systems.
@@
The third challenge is ensuring the safety and security of agentic applications. LLM agents, built using LLMs, are vulnerable to specific attacks such as prompt injection [17] and jailbreaking [18, 19]. Furthermore, in a global agentic framework, the use of dynamic capabilities can increase their exposure to widespread risks. For example, when LLM agents use third party APIs[20], malicious APIs can plan attacks and insert, substitute or delete key information. Depending on the agent’s application domain, such attacks could threaten physical security, financial security, or overall system integrity. This underscores the critical importance of applying robust safety and security measures in such frameworks.
@@
To address these three issues, the Distributed Agents in a Worldwide Network (DAWN) framework is proposed. DAWN’s architecture includes key components such as the Principal Agent and Gateway Agents, which work together to enable agents to plan, compose, discover, connect, orchestrate, observe, safeguard, and communicate globally. Given a user request, the Principal Agent autonomously plans, produces, submits a list of tasks describing the derived capabilities to Gateway Agents,and then orchestrates these resources to execute the plan. (A Principal Agent is connected with a fixed of Gateway Agents in a fashion similar to sub-contractors associated with a principal contractor on a manufacturing contract.) Gateway Agents may be distributed globally and have the capability to register public and proprietary resources (tools, agents, and agentic applications). Upon request, the Gateway Agents search their registries for the most suitable resources to address the proffered task list. These selected resources are wrapped into callable sequences and returned to the Principal Agent. Fig. 1 illustrates the globally distributed nature of this collaboration through the Principal Agent and Gateway Agents.
@@
DAWN offers multiple operational modes (and combinations of them) to orchestrate resources to accommodate tasks that require different level of determinism, as depicted in Fig. 2.:
@@
Figure 2: DAWN offers levels of determinism and autonomy required for different modes of agentic operation
@@
  • No-LLM Mode: This mode is best suited for tasks that require high predictability and reliability, leveraging traditional algorithms and tools to ensure consistent, repeatable results. Here, human operators manually design the workflow without using the Principal Agent and use only non-LLM tools retrieved from the Gateway Agents to maximize determinism.
@@
  • Copilot Mode: In this mode, agents act as intelligent assistants, enhancing deterministic workflows with creative input. Human operators still design the workflow without using the Principal Agent. Tools, agents, and agentic applications retrieved from the Gateway Agents are employed in the intermediate steps.
@@
  • LLM Agent Mode: In this mode, agents operate autonomously, making real-time decisions based on complex reasoning. Here, the Principal Agent defines the workflow and retrieves the necessary resources from the Gateway Agents autonomously. This mode is particularly effective in for complex tasks where business logic is fuzzy and there is tolerance for stochastic results.
The modes of operation are demonstrated in Section IV of the paper, where a human resources use case is used to exemplify the DAWN framework. DAWN implements safety, security, and compliance as a foundational layer within the architecture. It requires that all Gateway Agents meet minimum requirements across these three dimensions in order to participate. The DAWN framework has several key functions.
@@
  • Multi-agent collaboration: DAWN leverages Gateway Agents to link agents globally. These Gateway Agents maintain registries that store agent profiles and intelligently match them to tasks. Thus, DAWN builds agentic applications by facilitating collaboration among agents, enabling them to discover, communicate, and coordinate across different systems and organizations. This set of Gateway Agents facilitates the execution of complex tasks by harnessing the individual expertise and collective intelligence of agents collaborating effectively.
@@
  • Flexibility: DAWN introduces a versatile approach that adapts to a wide range of task requirements, from deterministic processes to dynamic decision-making. This adaptability allows for the seamless integration of traditional software tools, LLMs, and LLM agents, ensuring that the most appropriate method is used in each context.
@@
  • Interoperability: DAWN is designed for seamless integration with a broad spectrum of tools, LLMs, and agents, but wrapping them as REST APIs in a unified format. It ensures interoperability across diverse ecosystems, which makes it adaptable to wider range of applications that require a blend of agentic processes and traditional software systems.
@@
  • Connectivity: DAWN uses an open protocol to interact and communicate between distributed components. The Principal Agents maintains a channel that allows new Gateway Agents to request connections. These connections are accepted after due vetting. Similarly, communication between the Principal Agent and Gateway Agents occurs over encrypted channels using HTTPS. A similar arrangement exists between Gateway Agents and agentic resources.
@@
  • Comprehensive Security and Compliance: The inclusion of a dedicated safety, security, and compliance Layer underscores the architecture’s commitment to maintaining high standards of data protection, user privacy, and regulatory compliance. This layer ensures that agentic operations within the framework adhere to stringent security protocols, protecting against unauthorized access and ensuring that the system meets industry-specific compliance requirements. Different organizations will demand varying levels of safety and security for handling agentic inputs and outputs and may specify these as a condition for participating in that organization’s implementation of the DAWN architecture.
@@ II Related Works
II-A Large Language Models
Large Language Models have marked one of the most significant developments in artificial intelligence.Originally designed for tasks like machine translation and chat completion, state-of-the-art LLMs now demonstrate remarkable capabilities across a wide array of applications, including mathematics, coding, image analysis, image and video generation, speech recognition, and question answering. These successes have inspired researchers to propose the ”scaling law”[21], suggesting that continuous scaling of LLMs could eventually lead to the development of Artificial General Intelligence (AGI)[22]. Despite these impressive strides, one of the most pressing challenges is hallucinations—generating factually incorrect or nonsensical outputs[23, 24]. This issue arises because the internal knowledge within LLMs can be both limited and outdated, leading to errors that undermine their reliability and applicability in critical applications.
II-B LLM Agents
@@
II-C Multi-agent Systems
@@
Figure 3: DAWN reference architecture.
@@ III DAWN Architecture Detail
The DAWN architecture is a modular framework designed to orchestrate collaboration of distributed agents. It supports agentic applications to be built with various operational modes, including No-LLM mode, Copilot mode, Agent mode, and combinations of them. It supports tasks ranging from deterministic workflows (e.g., job offers and background checks) to creative and autonomous processes (e.g., interview summaries and job description compliance remediation). Central to the Agent mode is the Principal Agent, which plans tasks and requests corresponding resources (tools, LLMs, and agents), and multiple Gateway Agents, each of which independently oversees resource retrieval, validation, and execution required for the tasks identified by the Principal Agent. Multiple agentic design patterns may be adopted to balance the performance and cost when operating in the Agent mode. DAWN uses a robust communication layer to facilitate seamless interaction between the Principal Agent, Gateway Agents, and registered agents distributed around the globe. A context layer maintains memory and task history, ensuring that the requisite information is exchanged efficiently via the communication layer. The safety, security, and compliance layer is responsible for the adherence of Principal and Gateway Agents to strict standards for the safe operation of agents during data and control interchange. Details of the architecture are shown in Fig. 3.
@@ III-A Principal Agent
The Principal Agent acts as an autonomous planner-composer and central orchestrator within the DAWN framework. Its primary functions include creating a plan to execute customer requests, searching local resources (in the off chance some resources are available, a low probability occurrence), requesting additional resources from the Gateway Agents (the high probability occurrence), assessing resources returned from Gateway Agents for best fit, and delegating subtasks’ execution to the corresponding resources. Compared with the peer-to-peer communication pattern proposed by IoA[36] and AutoGen[10], using the Principal Agent as the central planner and orchestrator ensures an orderly and structured workflow that is efficient, reproducible, and easier to trace and restart if prematurely stopped. The Principal Agent first plans, seeks resources that meet its plan, and composes the task sequence using the retrieved resources. Orchestration, or the execution of the agentic resources, then begins with the planning using the cached information gathered during the compose phase to implement the sequence of resources required to fulfull the user’s intent. Other key components of the Principal Agent are as follows.
@@
Reasoning, Planning, and Adaptive Decision-Making: The Principal Agent utilizes an LLM to understand the user’s intent, develop a plan, and decompose it into manageable subtasks. This feature is evident when DAWN operates in copilot or LLM mode. The Principal Agent’s planner uses ReAct [28], as the default reasoning strategy, which follows a cycle of reasoning, acting, and observing to actively move through subtasks. This dynamic approach enables the Principal Agent to adjust to complex and changing environments. For example, the user might ask the Principal Agent to plan a trip. The Principal Agent would first interpret the goal end-to-end and then create a plan with subtasks like booking a flight, reserving accommodation, and arranging local transport. If the flight turns out to be unavailable, the ReAct strategy allows the Principal Agent to search for alternative way such as train and bus, and modify the plan on the fly. Depending on the task requirements, users can select other reasoning strategies. For less dynamic tasks where computational cost is a priority, the Reasoning Without Observation (ReWOO) strategy [37] may be more suitable. ReWOO creates a comprehensive, end-to-end plan without adjusting it during execution. HuggingGPT[38] uses this strategy to process multi-modal data by calling machine learning models available on HuggingFace[39] and achieves good performance while minimizing LLM calls. Alternatively, the Tree-of-Thoughts (ToT) strategy [29] encourages the agent to explore multiple reasoning paths and self-evaluate them, offering more deliberate decision-making but at a higher computational cost. This approach is beneficial when performance is critical and cost is secondary. In Language Agent Tree Search (LATS)[40] and ToolLLM[41], the authors combine the ToT strategy with other techniques to achieve the best success rate that surpasses ReAct.
Local Resource Pool: The Principal Agent hosts a local pool for storing and caching resources retrieved from gateway Agent. Multiple common resources, such as a calculation tool and a web search tool, are permanently held in the local resource pool. References to resources retrieved from the Gateway Agents are also cached in the local resource pool using a Least Recent Used (LRU) strategy. This design Principal helps the Principal Agent to respond to simple or repetitive (sub)tasks without repeatedly reaching out to Gateway Agents and dramatically reduces the communication overhead. This resource pool plays a critical role in the orchestration phase of the task sequence.
@@
Resource Requesting: After developing a plan, the Principal Agent summarizes the plan and its context—such as completed steps, conversation history, and user preferences—into queries and sends them to all connected Gateway Agents in an effort to retrieve the most appropriate resources to execute the plan in its entirety. If there are any suitable local resources, then the Principal Agent will not seek send out request for those.
@@
Resource Execution: Gateway Agents are expected to apply their programmed discretion in searching, matching and retrieving the most applicable resources. When requests are sent to multiple Gateway Agents, it is only natural that duplicate resources or resources with overlapping capabilities are returned to the Principal Agent. The Principal Agent is responsible for analyzing the returned resource manifest files, assessing the quality of each resource, determining which agentic resource or resources may be applied to each task list step, and then retaining just those resources. When the Planning Agent begins execution of the task graph, it continues until it hits a node where a new search for resources is required to complete the remaining subtasks. (This situation can arise if during the compose phase of the process, the returned resources from the Gateway Agents only partially fulfill the customer need..) The Planning Agent then reaches out to the Gateway Agents again. This process repeats until the user’s request is fully addressed. It should be noted that Gateway Agents return references to agentic resources. For instance, some agents may actually execute on their preferred cloud service provider locations. In other cases, the agents may rely on the run-time environment of their host Gateway Agents. In both cases, only callable references to resources are passed between agents, Gateway Agents and Principal Agents. And so, when the Principal Agents executes resources from a Gateway Agent, it is in reality making an API call to the resource and passing the payload to the concerned Gateway Agent via JSON or similarly packaged property file format.

### v3

@@ Abstract
The rapid evolution of Large Language Models (LLMs) has transformed them from basic conversational tools into sophisticated entities capable of complex reasoning and decision-making. These advancements have led to the development of specialized LLM-based agents designed for diverse tasks such as coding and web browsing. As these agents become more capable, the need for a robust framework that facilitates global agentic communication and collaboration for building sophisticated software solutions has become increasingly important. Distributed Agents in a Worldwide Network (DAWN) addresses this need by providing an architectural framework that allows globally distributed agents of any provenance to be registered, discovered, and organized for building AI-based applications and solutions. In DAWN, a Principal Agent Service composes and oversees the execution of agentic applications. It delegates tasks to one or more Gateway Agent Services that provide for the discovery, registration, and connection of the most suitable agents to fit each application’s need. DAWN offers three operational modes: No-LLM mode for deterministic and classical software development, Copilot for decision-making augmented using AI, and LLM Agent for autonomous operations. Last but not the least, DAWN ensures the safety and security of agent collaborations globally through a dedicated safety, security, and compliance layer, protecting the network against attackers and adhering to stringent security and compliance standards. These features make DAWN a robust framework for designing, developing and deploying agent-based applications across business and consumer applications.
@@ I Introduction
Rapid advancements in large language models (LLMs)[1, 2, 3] have have sparked significant interest in LLM-based AI agents[4] which have the potential to free humans from repetitive tasks and significantly enhance productivity. New (AI) agents are being developed on a regular basis and their performance is correspondingly improving. Agents exist for tasks such as coding[5, 6], web browsing[7], gaming, autonomous assistants, computer use, and others [8, 9]. Furthermore, multi-agent systems[10, 11, 12, 13, 14, 15, 16] have been proposed to leverage the collective intelligence and specialized profiles and skills of multiple agents. And agents are gradually finding their way into business processes and software development. However, three major challenges arise that hinder their potential for broader and more effective application.
@@
The first challenge is how to discover and connect to agents to accomplish tasks when they are distributed globally. Existing frameworks, such as LangGraph[14] and AutoGen[10] have bounded scope as they are typically designed to manage a limited number of agents that work well when created within their frameworks. This constraint arises from the need to manually hard-code the names, URLs, and description of other agents into the prompts, which becomes increasingly difficult as the network of agents grows. More agents created by different entities allow for a broader selection covering functional and non-functional capabilities. Additionally, integrating heterogeneous agents from different ecosystems requires developers to invest effort in developing wrappers and packages specific to each new framework as they evolve. New frameworks such as Agntcy[17], A2A[18], and ACP[19] mitigate some of the issues relating to finding and connecting agents.
@@
The second challenge arises because many business use cases demand determinism, consistency, and reliability, whereas LLM-based agents are inherently stochastic and excel in more flexible environments. They struggle with tasks requiring strict adherence to predefined rules and execution graphs. Empirical evidence shows that manually designing workflows, rather than granting full autonomy, leads to more robust performance and lower computational costs [20]. Balancing these often conflicting needs for autonomy and determinism is therefore essential to the broader adoption of agent-based systems. Moreover, agentic software development is expected to be a combination of AI and non-AI tools working together complicating the needs of business processes predictability while exploiting the power of AI.
@@
The third challenge is ensuring the safety and security of agentic applications. LLM-based Agents are vulnerable to specific attacks such as prompt injection [21] and jailbreaking [22, 23]. In a global agentic framework, the use of dynamic capabilities can increase their exposure to widespread risks. For example, when LLM agents use third party APIs[24], malicious APIs can plan attacks and insert, substitute, or delete key information. Depending on the agent’s application domain, such attacks could threaten physical security, financial security, or overall system integrity. This underscores the critical importance of applying robust safety and security measures in such frameworks. Add to this the fact that generalized LLMs are invoked via provider APIs external to most business and consumer networks, and data is pushed to them for analysis. This increases the risk of data exfiltration and concomitant privacy risks. Safety, security, and privacy of business and consumer data needs to be handled with great care. This requires the addition of control points during the design, build, and deploy phases of agentic software development.
@@
The proposed Distributed Agents in a Worldwide Network (DAWN) framework addresses these three issues. DAWN’s architecture includes key components such as the Principal Services Agent and Gateway Services Agent - Principal Agent and Gateway Agent, for short. These ensemble services work together to enable globally distributed agents to plan, compose, discover, connect, orchestrate, observe, safeguard, and communicate. Principal Agents are connected with a fixed set of Gateway Agents in a fashion similar to sub-contractors associated with a principal contractor on a manufacturing contract. Given a user request, the Principal Agent autonomously plans and then produces a list of tasks and agentic workflow describing the derived capabilities to subsidiary Gateway Agents. Gateway Agents, which may be local or distributed globally, have the capability to register public and proprietary resources (tools, agents, and agentic applications). Upon request, the Gateway Agents search their registries for the most suitable resources to address the proffered task list. These selected resources are wrapped into RESTful endpoints and returned to the Principal Agent. The Principal Agent then orchestrates these resources across Gateway Agents to execute the plan. Fig. 1 illustrates the globally distributed nature of this collaboration of an Principal Agent and Gateway Agents. As evident from the figure, the Principal Agent and attached Gateway Agents are more than mere single-function AI agents or applications. They host multiple functionalities and capabilities to support the creation, development, deployment, and operation of many applications that may be devised statically (predesigned and predeployed compute resources to address well-known workplace and consumer problems) or dynamically (on-the-fly accumulation and threading of AI agents and related resources along with the requisite compute infrastructure).
@@
DAWN offers multiple operational modes application designers can choose from to orchestrate resources to accommodate tasks that require different level of determinism, as depicted in Figure 2. Most new applications and solutions will likely integrate agents into existing business processes in copilot or autonomous mode as required.
@@
Figure 2: DAWN offers levels of determinism and autonomy required for different modes of agentic operation.
@@
  • No-LLM Mode: This mode is best suited for tasks that require high predictability and reliability, leveraging traditional algorithms and software tools to ensure consistent, repeatable results. Here, human operators manually design the workflow without using the Principal Agent and use only non-LLM resources and tools retrieved from the Gateway Agents to maximize determinism. In many ways, this is classical distributed software design and development.
@@
  • Copilot Mode: In this mode, agents act as intelligent assistants, enhancing deterministic workflows with creative input. Human operators design the workflow with the assistance of the Principal Agent’s LLM-based planner and composer while using many of its constituent services and capabilities. Humans in this mode serve as overseers of the task planners and agentic workflow composers. Tools, agents, and agentic applications retrieved from the Gateway Agents are employed in the intervening steps to complete and execute the agentic operations. This mode is a hybrid between classical distributed software development and AI-based agentic software development.
@@
  • LLM Agent Mode: In this mode, agents operate autonomously, making decisions based on complex reasoning driven by prompting, LLM-as-a-judge [25], fine-tuning, etc. Here, the Principal Agent defines the workflow and retrieves the necessary resources from the Gateway Agents autonomously. This mode is particularly effective in for complex tasks where business logic is fuzzy and there is tolerance for stochastic results. Software developers play a significant role in the application design, workflow development, and testing phases for business and consumer applications that demand robustness and repeatability. They may play a lighter role for applications that are conversational in nature.
DAWN implements safety, security, and compliance as a foundational layer within the architecture. It requires that all Gateway Agents meet minimum requirements across these three modes in order to participate. Principal Agents are always under the purview of the owning organization and hence all of the parent organization’s regulations apply to its construction and management, while Gateway Agents may be under the purview of partners, vendors, and other third parties. This means its internals may be proprietary and not transparent, but its capabilities, access methods, and communication channels are subject to conditions and constraints. The DAWN framework has several key functions.
@@
  • Multi-agent collaboration: DAWN builds agentic applications by facilitating collaboration among agents, enabling them to discover, communicate, and coordinate across different systems and organizations. Where more than one Gateway Agent is required to fulfill a set of tasks, the collected set of Gateway Agents facilitates the execution of complex tasks by harnessing the individual expertise of agents to collaborate effectively. DAWN leverages Gateway Agents to link agents globally. These Gateway Agents maintain agents themselves or link to registries that store agents’ profiles and destinations. In the event agent registries have tagged agents for use with tasks or if agent profiles carry track records of prior use, the Gateway Agent uses that information for agent selection. It may also use its intrinsic capabilities to perform additional inspection of agents in response to requests from the Principal Agent.
@@
  • Flexibility: DAWN introduces a versatile approach that adapts to a wide range of task requirements, from fully deterministic processes to developer-driven design to dynamic decision-making. This adaptability allows for the seamless integration of traditional software tools, LLMs, AI agents, and multi-agent systems ensuring that the most appropriate method is used in each context.
@@
  • Interoperability: DAWN is designed for seamless integration with a broad spectrum of software tools, LLMs, and AI agents, by wrapping them as REST APIs in a unified format. It ensures interoperability across diverse ecosystems, which makes it adaptable to wider range of applications that require a blend of agentic processes and traditional software systems.
@@
  • Connectivity: DAWN uses an open protocol to interact and communicate between distributed components. The Principal Agent maintains a separate channel that allows new Gateway Agents to request connections. These connections are accepted after due vetting. Similarly, communication between the Principal Agent and Gateway Agents occurs over encrypted channels using HTTPS. A similar arrangement exists between Gateway Agents and all agentic resources. Gateway Agents belonging to private entities may define their own private protocols and encrypted transport mechanisms to communicate with agent registries.
@@
  • Comprehensive Security and Compliance: The inclusion of a dedicated safety, security, and compliance layer underscores the architecture’s commitment to maintaining high standards of data protection, user privacy, and regulatory compliance. This layer ensures that agentic operations within the framework adhere to stringent security protocols, protecting against unauthorized access and ensuring that the system meets industry-specific compliance requirements. Different organizations will demand varying levels of safety and security for handling agentic inputs and outputs and may specify these as a condition for participating in that organization’s implementation of the DAWN architecture. These may be mandated by their own internal security, privacy, and compliance requirements, in additional to governmental regulations.
@@ II Related Works
II-A LLM Agents
@@
II-B Multi-agent Systems
@@
Figure 3: DAWN reference architecture
@@ III DAWN Architecture Detail
The DAWN architecture is a modular framework designed to orchestrate collaboration of distributed agents. It supports agentic applications to be built with various operational modes, including No-LLM mode, Copilot mode, and Agent mode, to support tasks ranging from deterministic workflows (e.g., job offers and background checks) to creative and autonomous processes (e.g., interview summaries and job description compliance remediation). Central to the Agent mode is the Principal Agent, which plans tasks and requests corresponding resources (tools, LLMs, and agents), and multiple Gateway Agents, each of which independently oversees resource retrieval, validation, and execution required for the tasks identified by the Principal Agent. DAWN uses a robust communication layer to facilitate seamless interaction between the Principal Agent, Gateway Agents, and registered agents distributed around the globe. A context layer within Gateway Agents maintains memory and task history, ensuring that the requisite information is exchanged efficiently via the communication layer. The safety, security, and compliance (SSC) layer is responsible for the adherence of Principal and Gateway Agents to strict standards for the safe operation of agents during data and control interchange. Details of the reference architecture are shown in Fig. 3.
Each organization may build and host one Principal Agent as the central portal for all of its users. From the organization’s software developer’s perspective, a desktop client is installed locally and includes access to their Principal Agent with its context, orchestration layer, and SSC layers, and one or more Gateway Agents. Some of these Gateway Agents may be entirely internal to the organization while others may be external with customizations to fit the organization’s particular needs. A developer may begin by selecting the appropriate operating mode for the application in question. If the user chooses the LLM Agent Mode, a chat-assistant interface is used to submit requests to the Principal Agent. If the application requested is a entirely new, the Principal Agent clarifies requirements with the user, interprets the request, devises a plan, and retrieves resources from Gateway Agents. (If a previously developed agentic workflow for that user task exists, then the Principal Agent presents that instead.) It then compiles an execution graph showing each subtask and corresponding resource, local or remote, which the user can approve. Upon approval, the orchestration layer oversees execution. Once execution completes, the user receives the final result. Because the Principal Agent runs locally in an organization’s network (or equivalently, in that’s organization virtual private cloud or VPC), user data resides in the local context layer, and transactions are overseen by the SSC layer, data privacy concerns are minimized.
@@ III-A Principal Agent
The Principal Agent acts as an autonomous planner-composer and central orchestrator within the DAWN framework. Its primary functions include creating a plan to execute customer requests, searching local resources, requesting additional resources from the Gateway Agents when necessary, assessing resources returned from Gateway Agents for best fit, and delegating subtasks’ execution to the corresponding resources. Compared with the peer-to-peer communication pattern proposed by IoA[36] and AutoGen[10], using the Principal Agent as the central planner and orchestrator ensures an orderly and structured workflow that is efficient, reproducible, and easier to trace and restart if prematurely stopped. The Principal Agent first plans, seeks resources that meet its plan, and composes the task sequence using the retrieved resources. Orchestration, or the execution of the agentic resources, then begins with the planning using the cached information gathered during the compose phase to implement the sequence of resources required to fulfill the user’s intent.
The Principal Agent is the primary interface within any organization to develop, deploy, and operate agentic applications and resources. End users may use applications already developed for most of their daily needs or develop new ones on the fly by assembling available agents. Organizations use software applications and tools developed by large software houses to run their businesses. In the agentic era, these will be replaced or augmented by agentic applications living alongside classical ones. Agents that fulfill many new software functions will become available via Gateway Agents. Gateway Agents serve as subcontractors for fulfilling agentic needs. The separation from the Principal Agent provides for separation of duties and function, private ownership, intellectual property protection, and varied levels of security and compliance flavors. Gateway Agents are by definition headless in the DAWN architecture. They exist to provide specialized and custom services.
The Principal Agent is a multi-tenant, multi-user application made of highly available constituent services that may be independently hosted and deployed across the organization, its private data centers, or on cloud data centers. In this regard, the Principal Agent is no different from an organization central web portal offering a host of services. The main difference is that the Principal Agent is also an autonomous AI-driven developer toolkit that can concurrently plan and compose multiple new applications and manage existing ones. And it has a built-in run-time environment to deploy and execute software. Each of its component services may be scaled independently. Principal Agents may be configured to reach out to internal and external Gateway Agents based on demand. Since most agents are hosted on or available via Gateway Agents, the availability of Gateway Agents serves as another scaling mechanism. Frameworks like Agntcy [17] envision globally replicated and synchronized agent directories (registries that describe agent function and capability). Gateway Agents will have ready access to these directories and registered agents on-demand.
@@
Reasoning, Planning, and Adaptive Decision-Making: The Principal Agent utilizes an LLM to understand the user’s intent, develop a plan, and decompose it into manageable subtasks. This feature is evident when DAWN operates in copilot or agent modes. The Principal Agent’s planner uses ReAct [29], as the default reasoning strategy, which follows a cycle of reasoning, acting, and observing to actively move through subtasks. This dynamic approach enables the Principal Agent to adjust to complex and dynamic environments. For example, an end-user might ask the Principal Agent to plan a trip. The Principal Agent would first interpret the goal end-to-end, and then create a plan with subtasks like booking a flight, reserving accommodation, and arranging local transport. If the flight turns out to be unavailable, the Principal Agent will search for alternative way such as train and bus, and modify the plan on the fly. The developer of the application decides how to prompt the planner component during development of this travel application. While ReAct works well in most use cases, developers may adopt other reasoning strategies. The Reasoning Without Observation (ReWOO) strategy [37] is more suitable for less dynamic tasks where computational cost is a priority. ReWOO creates a comprehensive end-to-end plan without adjusting it during execution. HuggingGPT[38] uses this strategy to process multi-modal data by calling machine learning models available on HuggingFace[39] and achieves good performance with minimal LLM calls. Alternatively, the Tree-of-Thoughts (ToT) strategy [30] encourages the agent to explore multiple reasoning paths and self-evaluate them, offering more deliberate decision-making but at a higher computational cost. This approach is beneficial when performance is critical and cost is secondary. In Language Agent Tree Search (LATS)[40] and ToolLLM[41], the authors combine the ToT strategy with other techniques to achieve the best success rate that surpasses ReAct. Developers may experiment with many of these before deciding on the best course that meets all of their application needs.
Local Resource Pool: The Principal Agent maintains a local resource pool that stores essential tools like calculators and web searchers, along with caching (references to) resources retrieved from Gateway Agents using a Least Recently Used (LRU) strategy. This setup enables the Principal Agent to efficiently handle simple or repetitive tasks without repeatedly consulting Gateway Agents, significantly reducing discovery and communication overhead while enhancing performance during task orchestration. The Principal Agent does not cache the agent itself if it is registered with a Gateway Agent – only its location and capability are used by reference within the application workflow. Security and compliance are example services whose agents may live within the Principal Agent’s orbit since they have a pivotal role to play in what information is sent to an LLM or received via chat response. Other resources that may exist locally include vector databases used in support of retrieval augmented generation (RAG) systems.
@@
Resource Requesting: After developing a plan, the Principal Agent summarizes the plan and its context – such as completed steps, conversation history, and user preferences – into queries and sends them to all connected Gateway Agents in an effort to retrieve the most appropriate resources to execute the plan in its entirety. Gateway Agents may not being able to fulfill the entire plan because they don’t have the resources to fulfill them all or because the resources required are down at the moment of request or because the task is niche and no agent exists for it. If there are any suitable local resources, then the Principal Agent will not seek send out request for those. With use, the Principal Agent’s contextual memory component learns both the functional and non-functional capabilities of different Gateway Agents it interacts with. Over time therefore the Principal Agent may acquire the necessary statistics to prefer some Gateway Agents over others for a given resource or capability. At the very least this gathered data can advise the agentic developer in the selection of preferred Gateway Agents in the same way sub-contractors earn the trust of the primary contractors in manufacturing or supply chains.
@@
Resource Execution: Gateway Agents are expected to apply their programmed discretion in searching, matching, and retrieving the most applicable resources. When requests are sent to multiple Gateway Agents, it is only natural that duplicate resources or resources with overlapping capabilities are returned to the Principal Agent. The Principal Agent is responsible for analyzing the returned resources, assessing the quality of each resource, determining which agentic resource or resources may be applied to each task, and then retaining just those resources. During the execution of the application workflow, the run-time component of the Principal Agent begins execution of the task graph. The execution continues until it hits a node where a new search for resources is required to complete the remaining subtasks. (This situation can arise if during the compose phase of the process, the returned resources from the Gateway Agents only partially fulfill the customer need.) The planner then reaches out to the Gateway Agents again with the remaining requirement. This process repeats until the user’s request is fully addressed. (In some cases, the response may be null and therefore some organizations may be better served if only fully qualified software applications with available agents are released to users.) It should be noted that Gateway Agents return references to agentic resources. For instance, some agents may actually execute on their preferred cloud service provider locations. In other cases, the agents may be imported into the Gateway Agent and rely on run-time environment of their host Gateway Agent. In both cases, only callable references to resources are passed between agents, Gateway Agents, and Principal Agents. And so, when the Principal Agents executes resources from a Gateway Agent, it is in reality making an API call to the resource and passing the payload to the concerned Gateway Agent via JSON or similarly packaged property file format.
In many agentic implementations, the action space or tool-set is limited to the local resource repository. This is typically due to three factors: 1) specialized agents naturally operate within a limited action space[42, 9, 43], 2) the context window of LLMs restricts the number of resources they can handle, and 3) if the agent has more than 20 tools to choose from, the accuracy in selecting the right tool drops as suggested by OpenAI[44]. (During experimentation the authors found this number to be less than 10.) The Model Context Protocol, MCP [45], was invented to address this shortcoming. However, fully autonomous agents need to navigate real-world problems where the action space is vast, and numerous resources are available. DAWN’s rationale to use distributed Gateway Agents follows directly from this tenet.